# Data-driven discovery of self-similarity using neural networks

### v1

@@ II.1 Dimension function
$U_{i}\mapsto U_{i}/L_{i}\quad\text{for each {\hbox{i}}}.$ (3)
The unit of $z_{i}$ is transformed under Eq. (3) as
@@
The function $\phi_{i}$ is called the dimension function. One can show that $\phi_{i}(\bm{L})$ is always written as a power-law monomial,222 This follows from the following argument. Let us consider the transformation by $\phi_{i}(\bm{L})$, $z_{i}\mapsto\phi_{i}(\bm{L})z_{i}$. This should be equivalent to the transformation by $\phi_{i}(\bm{L}^{\prime})$ followed by the transformation by $\phi_{i}\left(\frac{L_{1}}{L^{\prime}_{1}},\cdots,\frac{L_{N_{\rm unit}}}{L^{\prime}_{N_{\rm unit}}}\right)$. Thus, the function $\phi_{i}$ should satisfy the following consistency condition, $\frac{\phi_{i}(\bm{L})}{\phi_{i}(\bm{L}^{\prime})}=\phi_{i}\left(\frac{L_{1}}{L^{\prime}_{1}},\cdots,\frac{L_{N_{\rm unit}}}{L^{\prime}_{N_{\rm unit}}}\right).$ (5) Solutions of this equation are power-law monomials like Eq. (6).
@@ II.3 Scale transformations and data collapse
As a particular kind of scale transformations, let us introduce the scale transformation of units (e.g., $1\,{\rm m}\mapsto 1\,{\rm km}$), under which physical parameters change by a factor given by the dimension function,
@@
$\partial_{t}u(t,x)=D\partial_{x}^{2}u(t,x)\,,\quad u(0,x)=Q\delta(x)\,,$ (23)
where $Q$ is the total mass. In this case, the set of physical parameters are $\bm{z}=(u,t,x,D,Q)^{\top}$. We can construct two independent dimensionless parameters, $\pi_{1}=\frac{u\sqrt{Dt}}{Q}$ and $\pi_{2}=\frac{x}{\sqrt{Dt}}$. The self-similar solution of diffusion equation is in the form of $\pi_{1}=f\left(\pi_{2}\right)$. If we rescale the units of the length, mass and time as
$(U_{\rm length},U_{\rm mass},U_{\rm time})\mapsto\left(\frac{U_{\rm length}}{L},\frac{U_{\rm mass}}{M},\frac{U_{\rm time}}{T}\right)\,,$ (24)
the dimension function $\phi$ acting on $\bm{z}$ is given as
@@
Since $\pi_{1}$ and $\pi_{2}$ are dimensionless, they are invariant under $\phi$. Therefore, the transformation $\phi$ preserves the relation $\pi_{1}=f\left(\pi_{2}\right)$.
@@
When the set of scale transformations preserving $\bm{F}(\bm{z})=\bm{0}$ is identical to the scale transformations of units, i.e., $G^{\rm(s)}=G^{\rm(s)}_{\rm unit}$, the system only has self-similar solutions of the first kind.
@@
where $\bm{A}$ are parameters. By repeating the same argument with the dimension function (see Footnote 2), one can show that $\varphi_{i}$ is always written as a power-law monomial,
@@
Since the action of $G^{\rm(s)}_{\rm second}$ on $\bm{\pi}$ is given by Eq. (29), $Z_{i}$ is given by products some powers of dimensionless parameters. The construction of these parameters can be done in a completely similar manner to the introduction of $\bm{\pi}$. Namely, we can separate dimensionless parameters $\bm{\pi}$ into to groups,
@@
and one can introduce invariant parameters under $G^{\rm(s)}$ by rescaling $\pi^{({\rm II})}_{i}$ by a product of powers of $\pi^{({\rm I})}_{i}$,
@@
The function $\Phi(\bm{\pi})$ can be written in the form
$\Phi_{k}(\bm{\pi})=\prod_{i}(\pi^{({\rm I})}_{i})^{r_{ki}}\bar{\Phi}_{k}(\bm{Z})$ (35)
@@
Note that there is a hierarchical structure (see Fig. 1). While $\bm{\pi}$ that are composed of $\bm{x}$ are invariant under the scale transformations of units $G^{\rm(s)}_{\rm unit}$, $\bm{Z}$ are composed of $\bm{\pi}$ and are invariant under all the scale transformations. We will refer to $\bm{\pi}$ and $\bm{Z}$ as the similarity parameters of the first kind and similarity parameters of the second kind, respectively.
Figure 1: Hierarchical structure of similarity parameters. Dimensionless parameters $\bm{\pi}=\left(\pi_{1},\cdots,\pi_{N_{\pi}}\right)$ are referred to as similarity parameters of the first kind, that are invariant under the scale transformations of units. The similarity parameters $\bm{Z}=\left(Z_{1},\cdots,Z_{\bar{N}}\right)$ of the second kind are invariant under the all the scale transformations where $\bar{N}\coloneqq N_{\pi}-N$.
@@ III.1 Outline of the strategy
As discussed in previous section, the identification of power exponents of similarity parameters is sufficient to exploit scale-invariance of the problems. By the dimensional analysis, one can find the self-similar solutions of the first kind. When the problems belong to similarity of the second kind, the corresponding power-law exponents reflect a nontrivial physical law. Our primary goal is to obtain the power exponents of similarity parameters of the second kind in a data-driven way. Let us outline the strategy of our method in the following.
@@
Next, we perform dimensional analysis and introduce dimensionless parameters $\bm{\pi}$. Then, we check if the system allows for self-similar solutions of the second kind, which means to identify the similarity parameters of the second kind. Traditionally, this step involves making assumptions based on a physical law that is suitable for the specific problem, which can lead to biases.
To determine the power exponents of similarity parameters of the second kind, the framework of crossover of scaling laws summarized in Ref. [26] is useful. In many physical problems, there are certain region of parameters in which the dependencies on some parameters are weak and can be neglected, which results in a simple scaling law. In such an ideal region there is a constant dimensionless number composed of the scaling law,
@@ IV.1 Setting
$\Pi\coloneqq\frac{\delta_{m}}{R},\,\,\,\,\kappa\coloneqq\frac{h}{R},\,\,\,\,\eta\coloneqq\frac{\rho v_{i}^{2}}{E},\,\,\,\,\theta\coloneqq\frac{\mu}{E^{1/2}\rho^{1/2}R}.$ (66)
According to Ref. [26], using the dimensionless parameters, Eq. (65) is expressed by the following self-similar solution as
@@
In this problem, $\Psi=\frac{\Pi^{3}}{\kappa\eta}$ is a fixed similarity parameter of the second kind. It is derived from the Chastel-Gondret-Mongruel solution [35, 36, 37], which is the scaling law in elastic impact regime. Here we assume that the CGM solution represents the intermediate asymptotics in ideal region where $Z\ll 1$. That is, the similarity parameter $\Psi=\frac{\Pi^{3}}{\kappa\eta}$ approaches a constant as $Z\to 0$. Here we call $\Psi$ the CGM number. On the other hand, the driving parameter which breaks the asymptotics is $\theta$. Therefore, the interfering similarity parameter should include $\theta$.
@@ IV.2 Training with synthetic data
    where $\alpha_{X}$ is a random variable sampled from the standard normal distribution $\mathcal{N}(0,1)$, and $r$ is an adjustable parameter, common to all $X$, that represents the strength of the noise,666 For a given $X$, $\delta X$ is defined by sampling a random parameter $\alpha_{X}$ from the normal distribution $\mathcal{N}(0,1)$. If the sampling of $\alpha_{X}$ is repeated many times while $X$ is fixed, a set of data $\{X+\delta X\}$ will be obtained, which is scattered around $X$. The coefficient of variation of this data is nothing but $r$. In this sense, the parameter $r$ represents the strength of the noise. and equals to coefficients of variation.
@@
(a) The parameter $p_{1}$ as a function of $r$. The true value is $p_{1}=0$.
(b) The parameter $p_{2}$ as a function of $r$. The true value is $p_{2}=-1/2$.
@@
(a) The parameter $p_{1}$ as a function of $r$. The true value is $p_{1}=0$.
(b) The parameter $p_{2}$ as a function of $r$. The true value is $p_{2}=-1/2$.
@@
(a) The parameter $p_{1}$ as a function of $r$. The true value is $p_{1}=0$.
(b) The parameter $p_{2}$ as a function of $r$. The true value is $p_{2}=-1/2$.
@@ V.1 Synthetic system
To make a meaningful comparison, we have to fix the ambiguities of exponents discussed in Sec. III. The raw values of the exponents obtained from the trained neural network are in general linear combinations of $\bm{p}^{(0)}$ and $\bm{p}^{(1)}$. We fix the ambiguities as follows. Let us denote the raw parameters by $\bm{w}^{(0,0)}$ and $\bm{w}^{(0,1)}$101010 Even if the training is perfect, these parameters are linear combinations of $\bm{p}^{(0)}$ and $\bm{p}^{(1)}$, $\bm{w}^{(0,0)}=c_{0}\bm{p}^{(0)}+c_{1}\bm{p}^{(1)},\bm{w}^{(0,1)}=c_{2}\bm{p}^{(0)}+c_{3}\bm{p}^{(1)}$.. We first add a vector $\propto\bm{w}^{(0,1)}$ to $\bm{w}^{(0,0)}$ in such a way that the third components of the total vector is zero. Then, we further scale the obtained vector so that its zeroth component becomes 1, and the resulting vector is denoted by $\bm{p}^{(0)}$. Similarly, we add a vector $\propto\bm{p}^{(0)}$ to $\bm{w}^{(0,1)}$ so that the first component of the total vector is zero, and then we scale it to make the zeroth component to unity. The operation described above is summarized as follows:
@@ VI Conclusion and discussion
We have tested the method using synthetic and experimental data. In the case where the scaling function is assumed to have single argument, we have shown that our method works effectively using numerically generated data. We also applied our method to actual experimental data and gained non-trivial insights into self-similarity structure of the system. The method can be naturally generalized to the situations where the scaling function has multiple arguments. We confirmed that our method also works well in the case of two arguments using numerically generated data.
@@
    Number of similarity parameters of the second kind must be assumed.
@@
Regarding the second limitation, our strategy requires an initial assumption about the number of combinations of the similarity parameters of the second kind. This necessitates a process of trial and error. For instance, if assuming two invariants proves insufficient, we may need to explore the possibility of additional invariants.
The third limitation is more fundamental. Our method yields specific numerical values for power exponents, which may not suit problems where these exponents are nontrivial functions of dimensionless parameters, which corresponds to self-similar solutions of the second kind, Type B [^13] E.g., the self-similar solution of the second kind for the porous medium equation considering the fissures in the rock in Ch.3 of Ref. [3]. . If it is possible to determine the parameter range where the power exponents are viewed as unchanging, the present method continues to be of value. We are of the opinion that this situation can potentially be addressed by adjusting the exponents using supplementary neural networks. The practicality of this approach remains to be evaluated.

### v3

@@ II.1 Dimension function
$U_{i}\mapsto U_{i}/L_{i}\quad\text{for each {\hbox{i}}},$ (3)
where $L_{i}\in\mathbb{R}_{>0}$. The unit of $z_{i}$ is transformed under Eq. (3) as
@@
The function $\phi_{i}$ is called the dimension function. One can show that $\phi_{i}(\bm{L})$ is always written as a power-law monomial,
@@
where $\gamma_{ij}$ are power exponents.
This follows from the following argument. Let us consider the scale transformation with scale factor ${\bm{L}}=\left(L_{1},\cdots,L_{N_{\rm unit}}\right)$ by $\phi_{i}(\bm{L})$, ${\rm unit}_{i}\mapsto\frac{1}{\phi_{i}(\bm{L})}{\rm unit}_{i}$. This should be equivalent to the transformation with another scale factor ${\bm{L}^{\prime}}=\left(L^{\prime}_{1},\cdots,L^{\prime}_{N_{\rm unit}}\right)$ by $\phi_{i}(\bm{L}^{\prime})$ followed by the transformation by $\phi_{i}\left(\frac{L_{1}}{L^{\prime}_{1}},\cdots,\frac{L_{N_{\rm unit}}}{L^{\prime}_{N_{\rm unit}}}\right)$. Thus, the function $\phi_{i}$ should satisfy the following consistency condition,
$\frac{\phi_{i}(\bm{L})}{\phi_{i}(\bm{L}^{\prime})}=\phi_{i}\left(\frac{L_{1}}{L^{\prime}_{1}},\cdots,\frac{L_{N_{\rm unit}}}{L^{\prime}_{N_{\rm unit}}}\right).$ (6)
Solutions of this functional equation are power-law monomials like Eq. (5) [2, 3]. [^3] This argument will be repeated for an element $\varphi_{i}$, scale transformation acting on dimensionless parameters. How to solve the functional equation (5) is written in Ch.1 of Ref. [3]
@@ II.3 Scale transformations and data collapse
As a particular kind of scale transformations, let us introduce the scale transformation of units [^4] While it is called scale transformation of units, it does not mean that it actually changes the scale of units (e.g. 1 m $\to$ 100 cm). It is the scale transformations acting on the physical parameters (e.g. 1 m $\to$ 100 m) but their scale functions, which determine the factor by which the parameters change following their similarity, are equal to the dimension functions. In general, dimension functions are considered to act on the numerical values of physical quantities (see p.16 of Ref. [3]). Here we intend to discuss the self-similarity of the physical parameters, the relation between the self-similarity of units and the self-similarity of physical parameters., under which physical parameters change by a factor given by the dimension function,
@@
$\partial_{t}u(t,x)=D\partial_{x}^{2}u(t,x)\,,$ (23)
with the following initial and boundary conditions
$u(0,x)=Q\delta(x),\quad u(t,\pm\infty)=0,$ (24)
where $Q$ is the total mass, $u$ is the concentration of substance, $x$ is the location, $t$ is the time, $D$ is the diffusion coefficient, and $\delta(x)$ is the Dirac delta function. The set of physical parameters for this system is $\bm{z}=(u,t,x,D,Q)^{\top}$. We can construct two independent dimensionless parameters, $\pi_{1}=\frac{u\sqrt{Dt}}{Q}$ and $\pi_{2}=\frac{x}{\sqrt{Dt}}$. The self-similar solution of diffusion equation is in the form of $\pi_{1}=f\left(\pi_{2}\right)$. The dimension function $\phi$, scale transformation of units, acting on $\bm{z}$ is given as
@@
It is easy to see $\pi_{1}$ and $\pi_{2}$ are invariant under $\phi$. Therefore, the transformation $\phi$ preserves the relation $\pi_{1}=f\left(\pi_{2}\right)$. $\phi$ can be estimated by seeing the numerical values of $\bm{z}$ by rescaling the units of the length, mass and time as
@@
When the set of scale transformations preserving $\bm{F}(\bm{z})=\bm{0}$ is identical to the scale transformations of units, i.e., $G^{\rm(s)}=G^{\rm(s)}_{\rm unit}$, the system only has self-similar solutions of the first kind [^5] Note that similarity of the first kind corresponds to the case in which physical parameters and units posses the same structure of self-similarity as $G^{\rm(s)}_{\rm unit}$ is equivalent to the dimension functions that preserve the similarity under the scaling units. This is the reason why, in some case, the self-similarity can be exploited by simply applying dimensional analysis even though dimensional analysis should be nothing but the procedure to obtain the invariants of the scale transformation of the units, not the physical parameters. These are the cases when the problems belong to the similarity of the first kind..
@@
where $\bm{A}$ are scale factors. By repeating the same argument with the dimension function (see Sec. II.1), one can show that $\varphi_{i}$ is always written as a power-law monomial,
@@
Since the action of $G^{\rm(s)}_{\rm second}$ on $\bm{\pi}$ is given by Eq. (30), $Z_{i}$ is given by products some powers of dimensionless parameters. The construction of these parameters can be done in a completely similar manner to the introduction of $\bm{\pi}$. Namely, we can separate dimensionless parameters $\bm{\pi}$ into two groups,
@@
and the component of ${\bm{\pi}}^{({\rm I})}$ must be functionally independent one another, which means that the component of ${\bm{\pi}}^{({\rm I})}$ cannot be rescaled by the power-law monomials composed of the other components of ${\bm{\pi}}^{({\rm I})}$ to obtain invariant parameters under $G^{\rm(s)}$ [^7] This argument is exactly the repartition of $z^{({\rm I})}$, the parameters with independent dimensions in Subsec. B. In a more general discussion, it corresponds to the functionally independent parameters on the scale transformation, with which the parameters cannot be rescaled by other parameters to obtain invariant parameters. The number of parameters that can be reduced is equivalent to the dimension of the orbit of the scale transformation acting on the parameters. See pp. 86-89 of Ref. [33].. One can introduce invariant parameters under $G^{\rm(s)}$ by rescaling $\pi^{({\rm II})}_{i}$ by a product of powers of $\pi^{({\rm I})}_{i}$,
@@
The function $\Phi(\bm{\pi})$ can be written in the form,
$\Phi_{k}(\bm{\pi})=\prod_{i}(\pi^{({\rm I})}_{i})^{r_{ki}}\bar{\Phi}_{k}(\bm{Z}).$ (36)
@@
Note that there is a hierarchical structure (see Fig. 1). Physical parameters $\bm{z}$ are composed of a product of a numerical number and a physical units, which is not invariant under scale transformation of units. While $\bm{\pi}$ that are composed of ${\bm{z}}$ are invariant under the scale transformations of units $G^{\rm(s)}_{\rm unit}$, $\bm{Z}$ are composed of $\bm{\pi}$ and are invariant under all the scale transformations. We will refer to $\bm{\pi}$ and $\bm{Z}$ as the similarity parameters of the first class and similarity parameters of the second class, respectively [26].
Figure 1: Hierarchical structure of similarity parameters. Physical parameters $\bm{z}=\left(z_{1},\cdots,z_{N}\right)$ are parameters that are composed of a product of a numerical number and a physical unit. They are not invariant under the scale transformation of units. $\bm{\pi}=\left(\pi_{1},\cdots,\pi_{N_{\pi}}\right)$ are referred to as similarity parameters of the first class, that are composed of physical parameters and are invariant under the scale transformations of units. $\bm{Z}=\left(Z_{1},\cdots,Z_{\bar{N}}\right)$ are the similarity parameters of the second class, that are composed of similarity parameters of the first class and are invariant under the all the scale transformations where $\bar{N}\coloneqq N_{\pi}-N$. The dashed lines indicate the parameters with independent dimensions of their group (group ${\rm I}$) and the solid lines indicate the remaining parameters (group ${\rm II}$).
@@ III.1 Outline of the strategy
As discussed in previous section, the identification of power exponents of similarity parameters is sufficient to exploit scale-invariance of the problems. By the dimensional analysis, one can find the self-similar solutions of the first kind. When the problems belong to similarity of the second kind, the corresponding power-law exponents reflect a nontrivial physical law. Our primary goal is to obtain the power exponents of similarity parameters of the second class in a data-driven way. Let us outline the strategy of our method in the following.
@@
Next, we perform dimensional analysis and introduce dimensionless parameters $\bm{\pi}$. Then, we check if the system allows for self-similar solutions of the second kind, which means to identify the similarity parameters of the second class. Traditionally, this step involves making assumptions based on a physical law that is suitable for the specific problem, which can lead to biases.
To determine the power exponents of similarity parameters of the second class, the framework of crossover of scaling laws summarized in Ref. [26] is useful. In many physical problems, there are certain region of parameters in which the dependencies on some parameters are weak and can be neglected, which results in a simple scaling law. In such an ideal region there is a constant dimensionless number composed of the scaling law,
@@ IV.1 Setting
$\Pi\coloneqq\frac{\delta_{m}}{R},\,\,\,\,\kappa\coloneqq\frac{h}{R},\,\,\,\,\eta\coloneqq\frac{\rho v_{i}^{2}}{E},\,\,\,\,\theta\coloneqq\frac{\mu}{E^{1/2}\rho^{1/2}R},$ (67)
which were obtained by using the following physical parameters with independent dimensions, ${\bm{z}}^{({\rm I})}=\left(\rho,R,E\right)$. According to Ref. [26], using the dimensionless parameters, Eq. (66) can be expressed by the following self-similar solution as
@@
In this problem, $\Psi=\frac{\Pi^{3}}{\kappa\eta}$ is a fixed similarity parameter of the second class. It is derived from the Chastel-Gondret-Mongruel solution [35, 36, 37], which is the scaling law in elastic impact regime. Here we assume that the CGM solution represents the intermediate asymptotics in ideal region where $Z\ll 1$. That is, the similarity parameter $\Psi=\frac{\Pi^{3}}{\kappa\eta}$ approaches a constant as $Z\to 0$. Here we call $\Psi$ the CGM number. On the other hand, the driving parameter which breaks the asymptotics is $\theta$. Therefore, the interfering similarity parameter should include $\theta$.
@@ IV.2 Training with synthetic data
    where $\alpha_{X}$ is a random variable sampled from the standard normal distribution $\mathcal{N}(0,1)$, and $r$ is an adjustable parameter, common to all $X$, that represents the strength of the noise,101010 For a given $X$, $\delta X$ is defined by sampling a random parameter $\alpha_{X}$ from the normal distribution $\mathcal{N}(0,1)$. If the sampling of $\alpha_{X}$ is repeated many times while $X$ is fixed, a set of data $\{X+\delta X\}$ will be obtained, which is scattered around $X$. The coefficient of variation of this data is nothing but $r$. In this sense, the parameter $r$ represents the strength of the noise. and equals to coefficients of variation.
@@
(a) Parameter $p_{1}$ as a function of $r$. The true value is $p_{1}=0$.
(b) Parameter $p_{2}$ as a function of $r$. The true value is $p_{2}=-1/2$.
@@
(a) Parameter $p_{1}$ as a function of $r$. The true value is $p_{1}=0$.
(b) Parameter $p_{2}$ as a function of $r$. The true value is $p_{2}=-1/2$.
@@
(a) Parameter $p_{1}$ as a function of $r$. The true value is $p_{1}=0$.
(b) Parameter $p_{2}$ as a function of $r$. The true value is $p_{2}=-1/2$.
@@ V.1 Synthetic system
To make a meaningful comparison, we have to fix the ambiguities of exponents discussed in Sec. III. The raw values of the exponents obtained from the trained neural network are in general linear combinations of $\bm{p}^{(0)}$ and $\bm{p}^{(1)}$. We fix the ambiguities as follows. Let us denote the raw parameters by $\bm{w}^{(0,0)}$ and $\bm{w}^{(0,1)}$141414 Even if the training is perfect, these parameters are linear combinations of $\bm{p}^{(0)}$ and $\bm{p}^{(1)}$, $\bm{w}^{(0,0)}=c_{0}\bm{p}^{(0)}+c_{1}\bm{p}^{(1)},\bm{w}^{(0,1)}=c_{2}\bm{p}^{(0)}+c_{3}\bm{p}^{(1)}$.. We first add a vector $\propto\bm{w}^{(0,1)}$ to $\bm{w}^{(0,0)}$ in such a way that the third components of the total vector is zero. Then, we further scale the obtained vector so that its zeroth component becomes 1, and the resulting vector is denoted by $\bm{p}^{(0)}$. Similarly, we add a vector $\propto\bm{p}^{(0)}$ to $\bm{w}^{(0,1)}$ so that the first component of the total vector is zero, and then we scale it to make the zeroth component to unity. The operation described above is summarized as follows:
@@ VI Conclusion and discussion
We have tested the method using synthetic and experimental data.161616 Additionally, we tested our method on data without self-similarity of the second kind. We revealed several indicators suggesting the absence of self-similarity in the data. See Appendix B for details. In the case where the scaling function is assumed to have single argument, we have shown that our method works effectively using numerically generated data. We also applied our method to actual experimental data and gained non-trivial insights into self-similarity structure of the system. The method can be naturally generalized to the situations where the scaling function has multiple arguments. We confirmed that our method also works well in the case of two arguments using numerically generated data.
@@
In certain cases, however, the data itself may exhibit hysteresis or bivalued behavior [6, 7], that can introduce instability into the estimation process. When such behavior is present, the accuracy of the estimation can be significantly improved by restricting the analysis to data points within the single-valued region, as discussed in Appendix A.
@@
    Number of similarity parameters of the second class must be assumed.
@@
Regarding the second limitation, our strategy requires an initial assumption about the number of combinations of the similarity parameters of the second class. This necessitates a process of trial and error. For instance, if assuming two invariants proves insufficient, we may need to explore the possibility of additional invariants. We find that the estimation becomes visibly unsatisfactory when we have a false assumption about the number of invariants (see Appendix B).
The third limitation is more fundamental. Our method yields specific numerical values for power exponents, which may not suit problems where these exponents are nontrivial functions of dimensionless parameters, which corresponds to self-similar solutions of the second kind, Type B [^18] E.g., the self-similar solution of the second kind for the porous medium equation considering the fissures in the rock in Chapter 3 of Ref. [3]. . If it is possible to determine the parameter range where the power exponents are viewed as unchanging, the present method continues to be of value. We are of the opinion that this situation can potentially be addressed by adjusting the exponents using supplementary neural networks. The practicality of this approach remains to be evaluated.
@@
The source codes and the data for this article are openly available from Ref. [24].